The weight of a diesel engine is directly proportional to the compression ratio (W=kr) because extra metal must be used to strengthen the engine for the higher pressures. Examine the net specific work produced by a diesel engine per unit of weight as the pressure ratio is varied and the specific hear input remains fixed. Do this for several heat inputs and proportionality constraints k. Are there any optimal combinations of k and specific heat inputs?
Answer #1

from the question. volume of the combustion chamber to the volume work from Tocto BDC & BPC to To PAGE 1 *Solution : ConsidelPAGE-2 Efficiency na Compression Ratio de Vive = // 1+ N=1t n = 14 Cv (To-T2) cpltto) n = 1-16 I és 041-5) VE P2 ECV (Tarta)PAGE 3 3 Wq3 .اونا Hele w2-3 ~ constant prosure expension Wzy Adiebotic expension W = Spoi = P(V3-V2) dan Bar = m.au = mc, AT